OpenStudy (mendicant_bias):

I'm dealing with some laplace transforms for the first time regarding trig functions mixed with Heaviside and am running into trouble; posted below momentarily.

OpenStudy (mendicant_bias):

\[\mathcal{L}\left\{\cos(2t)\mathcal{U}(t-\pi)\right\}\]

OpenStudy (mendicant_bias):

I'm confused about how to put the argument of the cosine function in terms of t-pi; I'm thinking of using trigonometric identities, but I can't find any looking through pages that deal with pi, only pi/2.

OpenStudy (mendicant_bias):

@midhun.madhu1987

OpenStudy (mendicant_bias):

@Kainui :|

OpenStudy (midhun.madhu1987):

I am not aware of this :( So sorry

OpenStudy (mendicant_bias):

No problem, thank you nonetheless!

OpenStudy (kainui):

If you have something with pi/2 surely if you apply it twice you can get yourself to pi? As for the Heaviside function is concerned, all that really does is means that from -infinity to pi the function is 0 and from pi to +infinity the function is normal. So just adjust your bounds of integration accordingly and play around with your identities.

ganeshie8 (ganeshie8):

period of cos(2t) is pi so cos(2t) = cos(2t \(\pm\) pi)

OpenStudy (mendicant_bias):

I know how the Heaviside function works and how the lower limit contribution disappears in the derivation for it; I'm just confused about how I should play around with those identities. I'm not sure how to move forward from what @ganeshie8 posted, but I'm thinking over it.

ganeshie8 (ganeshie8):

does this help cos(2t) = cos(2t-2pi) = cos(2(t-pi))

OpenStudy (mendicant_bias):

Oh, yeah, hah, that makes sense, I understand that more than I do the first mention. I want to think over your first post for a minute, but I can see how the second one would work.

OpenStudy (mendicant_bias):

Alright, so from the two posts, I'm confused on two different accounts unless there was a typo in the first one. Changing the argument of a trig function changes its period, so-in agreement with the first post-\[\cos(2t)=\cos(2t \pm \pi)\] But in the second post, it looks like the periodicity of it didn't change despite being multiplied by 2, and 2pi is still being used for the period.

ganeshie8 (ganeshie8):

adding / subtracting ANY natural number of `pi`s from the argument of cos(2t) wont change its value cuz cos(2t) repeats itself after every cycle of period, pi.

ganeshie8 (ganeshie8):

cos(2t + pi) = cos(2t + pi + pi) = cos(2t + pi + pi + ... )

OpenStudy (mendicant_bias):

Oh, yeah, derp.

OpenStudy (mendicant_bias):

Alright, so \[\mathcal{L}\left\{\cos(2t)\mathcal{U}(t-\pi)\right\}=\mathcal{L}\left\{\cos(2t-2\pi)\mathcal{U}(t-\pi)\right\}=\mathcal{L}\left\{\cos(2(t-\pi)\mathcal{U}(t-\pi)\right\}\]

OpenStudy (mendicant_bias):

From there, the identification\[a=\pi, \ f(t)=\cos(2t)\]

OpenStudy (mendicant_bias):

\[\mathcal{L}\left\{\cos(2t)\mathcal{U}(t-\pi)\right\}=\frac{se^{-\pi s}}{s^2+4}\]

OpenStudy (mendicant_bias):

I'm not sure what you mean, sorry. I get it now, though.

ganeshie8 (ganeshie8):

http://prntscr.com/58i841

ganeshie8 (ganeshie8):

nvm i see what you're doing now :)

OpenStudy (mendicant_bias):

ty!